OpenStudy (timothygeorge04):

This might sound ridiculous, but it's been so long out of school. What's the proper way to do multiplication in your head, please explain. 12x 7.43 =

OpenStudy (photon336):

well it's going to be hard to do that in your head because of the decimals. best way would be to start with a multiplication table, and work your way up from there. for example let's start with 1 anything multiplied by 1 is just itself 1*1 = 1 2*1 = 2 180*1 = 2 if you multiply a number by 2 it's the same as if you took the same number and added it twice. for example 2*4 = 8 4+4 = 8 8*2 = 16 8+8 = 16

OpenStudy (photon336):

again lol typo's 180*1 = 180

OpenStudy (astrophysics):

Just the way you'd it on paper! That could work just as well!

OpenStudy (photon336):

If someone asked me that off the spot I wouldn't be able to give an exact number. i would round 7.43 to 7 and just say 7*12 which is easier.

OpenStudy (shadowlegendx):

If you are good at retaining images in your mind and painting on them  You could do it that way. Just mentally set it up then start multiplying